OpenStudy (anonymous):

If x^2+2=6x is solved by completeing the square, an intermediate step would be 1)(x+3)^2=7 2)(x-3)^2=7 3)(x-3)^2=11 4)(x-6)^2=34

OpenStudy (anonymous):

(x-3)^2=7

OpenStudy (cwrw238):

(x -3)^2 - 9 = -2 (x -3)^2 = 7

OpenStudy (anonymous):

what foolformath said

jhonyy9 (jhonyy9):

2. is right